Grade 11Physical Chemistry

Silver (108 gm/ mol) has a density of 10.5 gm/ cm^3 the no. Of silver atoms on a surface area 10^(-22)m^2 can be expressed as y×10^x the value of x is A.2. B.9 C.8 D.4

Tanay bhadula
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
d=mv  10.5 g/cc
Number of atoms of Ag in 1cc ⇒ 10.5108×NA
In 1cm, number of atoms of  Ag =  310.5108×NA−−−−−−−−√
In 1cm2
, number of atoms of Ag = =(10.5108×NA)2/3
In 10–12mor 10–8cm2
, number of atoms of
Ag=(10.5108×NA)23 × 10−8                             =((10.5×6.02×1023)108)23 × 10−8
=1.5×107
Thus, x = 7